OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -1/2x + 5. a. y = -2x + 2 b. y = 2x - 14 c. y =1/2x - 8 d. y =-1/2x - 4

OpenStudy (anonymous):

Your given line is in slope-intercept form: y = mx + b where "m" is the slope and that value is "-1/2". You want a line that is parallel to the given line, so you will have another line with the same slope. You could just stop at this point, because there is only one choice with a slope that is "-1/2". I could show you how to do this problem if you were given no choices if you like.

OpenStudy (anonymous):

From my first post, you already have the correct answer, but if you were given no choices, then you first have to use the point-slope form for the equation of a line:\[y - y _{1} = m(x - x _{1})\]where x1 and y1 are from your given point. y - (-6) = (-1/2)(x - 4) and then just simplify this equation into the slope-intercept form: y = mx + b

OpenStudy (anonymous):

All good now @jen1511 ?

OpenStudy (anonymous):

Yes hahah. thanks :)